OpenStudy (anonymous):

CHECK MY WORK! How many solutions will this system of equations have? y = −3.5x − 3.5 y = −3.5x + 3.5 No solution<<<<

OpenStudy (anonymous):

Two solutions, x = + 1, -1

OpenStudy (anonymous):

3.5 = -3.5x 3.5 / -3.5 = x x = -1

OpenStudy (anonymous):

There is no solution. They are parallel.

OpenStudy (anonymous):

They are parallel, they have the same slope. They just have different y-intercepts.

OpenStudy (anonymous):

which answer?

OpenStudy (anonymous):

The first one

OpenStudy (anonymous):

No, if they are parallel the set has infinite solutions.

OpenStudy (anonymous):

I suggest using geogebra.

OpenStudy (anonymous):

Solution means intersecting points.

OpenStudy (anonymous):

There are no intersecting points in parallel lines.

OpenStudy (anonymous):

Oh, crap. You're right.

OpenStudy (anonymous):

this is algebra 1. So no solution?

OpenStudy (anonymous):

I'm overthinking the entire thing.

OpenStudy (anonymous):

Yes, there is no solution.

OpenStudy (anonymous):

A dependent system of equations is a system with __________. the same line parallel lines intersecting lines and lines that have the same equation intersecting lines and lines that have the same slope

OpenStudy (anonymous):

Dependent system of equations are the same line.

OpenStudy (anonymous):

so A?

OpenStudy (anonymous):

Independent system of equations intersect one another, as opposed to dependent. And yes, it's A.

OpenStudy (anonymous):

Solve the system of equations. x + 3y = −1 2x + 2y = 6 (−4, 1) (2, −1) (4, −1) (5, −2) @izuru

OpenStudy (anonymous):

I previously answered a question similar to this, hold on.

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

Like, it's the exact same problem. I couldn't find it but I'll still help you out lol So you can use substitution in this one. Try and isolate x in the first equation.

OpenStudy (anonymous):

x+3y=-1 -3y -3y x=-4y

OpenStudy (anonymous):

-3y and -1 are not like terms.. So it's just x = -3y -1. Now, plug that isolated x into the second equation. 2x + 2y = 6 2(-3y - 1) + 2y = 6

OpenStudy (anonymous):

2(-1)+(-1y)=6

OpenStudy (anonymous):

You have to distribute 2*(-3y - 1) before adding +2y.

OpenStudy (anonymous):

2*-3y-2

OpenStudy (anonymous):

@izuru

OpenStudy (anonymous):

The equation turns into -2 -6y + 2y = 6. You multiply 2 by -3y and -1.

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

so B? @izuru

OpenStudy (anonymous):

@IrishBoy123 b?

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

@freckles @Hero

OpenStudy (anonymous):

@CandyCove

OpenStudy (anonymous):

b

OpenStudy (anonymous):

y = -2. Plug it in with the first equation to figure out the value of x.

OpenStudy (anonymous):

Its D! @izuru

OpenStudy (anonymous):

You're right